# Experimental Measurements of the Granular Density of Modes via Impact

### v1

@@ Abstract
The jamming transition is an important feature of granular materials, with prior work showing an excess of low frequency modes in the granular analog to the density of states, the granular density of modes. In this work, we present an experimental method for acoustically measuring the granular density of modes using a single impact event to excite vibrational modes in an experimental, three dimensional, granular material. We test three different granular materials, all of which are composed of spherical beads. The first two systems are monodisperse collections of either 6 mm or 8 mm diameter beads. The third system is a bidisperse mixture of the previous two bead sizes. During data collection, the particles are confined to a box; on top of this box, and resting on the granular material, is a light, rigid sheet onto which pressure can be applied to the system. To excite the material, a steel impactor ball is dropped on top of the system. The response of the granular material to the impact pulse is recorded by piezoelectric sensors buried throughout the material, and the density of modes is computed from the spectrum of the velocity autocorrelation of these sensors. Our measurements of the density of modes show more low frequency modes at low pressure, consistent with previous experimental and numerical results, as well as several low frequency peaks in the density of modes that shift with applied pressure. Finally, we also see that the density of modes falls off at a wavelength on the order of twice the particle diameter, which is reminiscent of the Debye frequency.
@@ I Introduction
Ranging from the rings of Saturn to a pile of sand on the beach, granular materials are ubiquitous throughout the Universe, yet many important granular phenomena are not fully understood. One such phenomena is jamming [1], in which a granular material can undergo a transition from solid-like to liquid-like behavior or vice versa. Jamming is an important phenomena for understanding not only the fundamental behavior of granular materials but also many industrial [16] and geophysical [5, 17] processes. For example, how close or far a material is from the jamming transition is studied in geophysics for the purposes of predicting soil stability and events such as landslides. Additionally, jamming also has application for the manufacturing of pharmaceuticals as much of this process involves the manipulation of granular materials [11].
One particularly interesting phenomena that is associated with the jamming transition is that an excess number of vibration modes (soft modes) are seen as the material approaches the jamming transition. As such, the density of states (also known as the density of modes when applied to granular materials) can be used to study the jamming transition, and potentially be used to characterize how far the material is from the jamming transition. Several previous simulations and experiments have done this for granular [8, 14, 18, 13, 4, 2] and colloidal systems [3, 6] using a variety of methods. However, many of the current methods used to measure the granular density of modes are cumbersome and not well suited to use outside of a laboratory setting. This work seeks to develop a simplified measurement process, particularly in regards to phonon excitation, that will aid in more widespread accessibility and deployment of density of modes measurements in non-laboratory, field conditions.
In this paper, we will experimentally measure the density of modes of a real three-dimensional granular material utilizing the spectrum of the velocity autocorrelation function. This is a standard method used in solid state physics [7] and has been experimentally applied to granular materials [13, 4, 2]. A brief summary of the method is as follows: first, the velocity autocorrelation function, $C_{v}(T)$, of a thermal system is given by
@@
where $v_{j}(t)$ is the velocity as a function of time of the $j$ particle, and the sum is over all the particles in the system. Then, the density of states, $D(\omega)$, is given by the Fourier transform of the velocity autocorrelation function
@@
In particular, when using the normalized autocorrelation function defined in Equation 1, the density of modes is normalized as follows:
$\int_{0}^{\omega_{D}}D(\omega)\mathrm{d}\omega=1$ (3)
where $\omega_{D}$ is the highest frequency mode in the system, and in the Debye model, $\omega_{D}$ has a wavelength that is twice the inter-atomic separation.
This method of finding the density of states assumes that the system is in thermal equilibrium and that the equipartition theorem holds. It also assumes access to the velocities of all the particles. However, even though granular materials are athermal systems, this method was applied to a 2D granular material by exciting the granular material with a white noise acoustic signal to mimic thermal motion [13]. This complex method of exciting the granular material proved successful and agreed well with previous simulations, even though it only utilized a subset of the particles. Building on this method, another experiment [4] used the natural phonon emissions of stick slip events of a granular material under shear to compute the density of modes. A key aspect of utilizing the spectrum of $C_{v}$ to find the granular $D(\omega)$ is that the granular material needs to be perturbed so that velocities can be measured.
In this paper, we present a novel approach to exciting the granular material so that the granular density of modes can be found using the spectrum of the velocity autocorrelation function. Our method excites the vibrational modes using a small impact on top of the granular material. This impact will create a pulse containing a wide spectrum of frequencies. The range of frequencies in a pulse is inversely proportional to the width of the pulse, with an infinitesimal pulse having a flat spectrum. This wide injection of frequencies will allow the granular material to vibrate more strongly at its natural resonances and attenuate signals that are not resonant with the natural modes of the system. This will effectively produce quasi-thermal vibrations among the grains.
This method is functionally very similar to the white noise excitation used in [13]; however, generating an impact is experimentally simple, leading to an improved technique that can be more broadly applied. Our results show that this technique yields a density of states that is consistent with the prior work in that the system has an excess of low frequency modes as the jamming transition is approached. Additionally, we are able to identify several peaks in the density of modes that shift with pressure (distance from the jamming transition).
@@
Figure 1: (a) A schematic of the experimental setup. The granular material, shown as the dotted fill, is confined inside a box, shown as the black line surrounding the granular material. Inside the granular material, labeled with a “P” in the figure are the piezoelectric sensors. On top of the granular material is the “Pressure Plate” assembly. The pressure plate assembly houses the force sensors that measure the pressure applied to the granular material via the “Screw” which is itself confined to the “Frame.” Vibrations are excited in the granular material when the “Ball” is released from the solenoid and dropped through a tube onto the bottom plate of the pressure plate assembly. The bottom plate of the pressure plate assembly is in direct contact with the top of the granular material. (b) A photo showing the ball dropping apparatus with the red wires composing the solenoid and the aluminum tube below the solenoid. To the right, and from the top down, are as follows: the piezoelectric sensor encased in the ABS half spheres, the bare piezoelectric sensor, a 6 mm grain, a 8 mm grain, and the steel impactor ball.
@@ II Experimental Methods
[…]
II.1 Phonon Excitation
Next, since the granular material is athermal and not in continuous motion, it must be excited in order to vibrate. In order to excite the granular material to vibrate at its natural modes, we drop a 1.05 g, 6.35 mm diameter, steel ball, see Figure 1, on top of the system, thereby injecting an acoustic pulse into the system. As described earlier, a pulse is used due to the fact that it contains a wide range of frequencies, and will produce vibrations in the granular material similar to thermal vibrations.
In order to produce a consistent pulse across our different packings and pressures, the following apparatus was constructed to drop the pulse-producing steel ball on the granular material. The pulse producing apparatus (labeled “ball” in Figure 1) consists of a 5.0 cm tall hollow aluminum tube on top of which is a solenoid. Inside and at the top of the tube, but below the solenoid, the steel ball is placed. The steel ball is held in place by the solenoid’s magnetic field and can be dropped when the solenoid is turned off. This setup allows for a consistent kinetic energy of the ball (0.5 mJ) as it strikes the top plate of the system, as well as provides a means of coordinating the time at which the ball is dropped with the data acquisition of the embedded piezoelectric sensors. Additionally, by striking the rigid plate of the pressure plate apparatus, the energy of the pulse is more evenly distributed across the granular material rather than being concentrated in a particular region of the system.
II.2 Piezoelectric Sensors
We use five piezoelectric sensors embedded in the granular material to measure the acoustic vibrations of the system caused by the steel ball being dropped on the granular material. The particular piezoelectric sensors used are Murata 7BB-12-9 buzzers. While these crystals are marketed as buzzers, they work equally well as detectors due to the symmetric nature of the piezoelectric effect: that is, a strain on a piezoelectric crystal will produce charge separation and, inversely, a charge difference across a piezoelectric crystal will produce a strain. The piezoelectric sensors are 12 mm in total diameter with the crystal being 9 mm in diameter. For each sensor, we glue two ABS plastic half spheres on either side in order to turn the sensor into a 12 mm diameter ABS sphere with a piezoelectric sensor running through its equator, see Figure 1. Encasing the sensors in ABS spheres allows the sensor to better blend in with the rest of the material which is also composed of ABS plastic spheres. This helps to ensure that the sensors are vibrating in as close a manner to the bulk granular material as possible.

### v2

@@ Abstract
The jamming transition is an important feature of granular materials, with prior work showing an excess of low frequency modes in the granular analog to the density of states, the granular density of modes. In this work, we present an experimental method for acoustically measuring the granular density of modes using a single impact event to excite vibrational modes in an experimental, three dimensional, granular material. We test three different granular materials, all of which are composed of spherical beads. The first two systems are monodisperse collections of either 6 mm or 8 mm diameter beads. The third system is a bidisperse mixture of the previous two bead sizes. During data collection, the particles are confined to a box; on top of this box, and resting on the granular material, is a light, rigid sheet onto which pressure can be applied to the system. To excite the material, a steel impactor ball is dropped on top of the system. The response of the granular material to the impact pulse is recorded by piezoelectric sensors buried throughout the material, and the density of modes is computed from the spectrum of the velocity autocorrelation of these sensors. Our measurements of the density of modes show more low frequency modes at low pressure, consistent with previous experimental and numerical results, as well as several low frequency peaks in the density of modes that shift with applied pressure. Our method represents an experimentally simple technique for investigating the granular density of modes and may increase the accessibility and number of such measurements.
@@ I Introduction
Jamming is the phenomena in which a granular material becomes rigid and able to support a finite shear stress and pressure [1, 2, 3, 4, 5, 6]. The jamming transition is important for understanding not only the fundamental behavior of granular materials but can also be applied in diverse industrial settings ranging from universal, soft, robotic grippers [7, 8, 9] to hopper flow [10, 11, 12] to the manufacture of pharmaceuticals [13]. Jamming also has application on a wide variety of geophysical processes [14, 15] such as earthquakes [16, 17, 18] and avalanches [19].
The jamming transition exhibits features of a second order phase transition; specifically, diverging length scales are observed as jamming is approached [20, 21, 22, 3, 23, 24]. In particular, simulations of the granular density of states [20, 3, 25, 26] observe long wavelength modes that diverge as the jamming transition is approached. These simulations of the density of states, $D(\omega)$, display an excess number of low frequency modes as compared to Debye scaling [27], $\omega^{d-1}$, where $d$ is the dimensionality of the system and $\omega$ is the frequency. These simulations identify the diverging length scale by first identifying a characteristic frequency, $\omega^{*}$, that is either the frequency at which the density begins to fall to zero at low frequency [20], or the frequency at which $D(\omega)$ falls to some fixed fraction of its maximum value [25] at low frequency. Regardless of how $\omega^{*}$ is defined, it is observed to have a power-law relationship with proximity to the jamming transition as quantified by excess coordination number $(z-z_{c})$ and relative packing fraction $(\phi-\phi_{c})$ [20, 25, 24].
In this work, we present a technique for experimentally measuring the granular density of states. We thus begin by briefly summarizing some of the previously used methods for measuring and calculating $D(\omega)$. Simulations calculate the density of states using the dynamical matrix constructed from the particle positions and grain interaction potential. A related method uses a covariance matrix; this method has been used to experimentally measure $D(\omega)$ in a two-dimensional (2D) colloidal systems [28, 29] by utilizing particle tracking to construct the covariance matrix from the particle displacements induced from thermal motion. However, applying this method to granular experiments requires an external vibration as a substitute for the thermal vibrations naturally present in colloids. Experiments in a 2D granular system [30] applied such an external vibration to their system and used particle tracking to construct a covariance matrix; these experiments found the number of long wavelengths mode increased as jamming was approached. While the previously mentioned matrix methods are successful in finding the vibrational modes of the system, these types of experimental methods are generally reserved for 2D systems that are small enough for the particles to be tracked. This presents an impractical option for large 3D systems. Another method employed in Refs. [31, 32] uses the spectrum of the velocity autocorrelation of the particles to measure the granular density of states (VACF method) and can be applied to large, 3D systems.
As the VACF method described in Ref. [31] will form the basis of our work, we will describe this method in detail. First, the velocity autocorrelation function, $C_{v}(t)$, is defined as follows
@@
where $v_{j}(t)$ is the velocity as a function of time of the $j$ oscillator in the system. $v_{j}(t)$ can be written as
$v_{j}(t)=A_{j}\omega_{j}\sin(\omega_{j}t+\phi_{j})$ (2)
where $A_{j}$ is the amplitude of the oscillation, $\omega_{j}$ is the angular frequency of the oscillator, and $\phi_{j}$ is the phase of the oscillator. The sum in Equation 1 is over all the oscillators in the system. We now briefly summarize the derivation presented by Ref. [33] in order to clearly identify the assumptions and conditions of this method. The first step of this derivation substitutes $v_{j}$, as defined in Equation 2, into $C_{v}(t)$, as defined in Equation 1. $C_{v}(t)$ is then simplified using the following assumptions: the first assumption is that the system is in a steady state such that $A_{j}$ does not vary with time. The second assumption is that the equipartition theorem is applicable to the system, meaning that each oscillator has an energy, on average, of $kT$, where $k$ is the Boltzmann constant and $T$ is the temperature. The third assumption is that the phase of the oscillators are random. Using these assumptions, $C_{v}(t)$, simplifies to
$C_{v}(t)=\frac{1}{N}\sum_{j}\cos(w_{j}t)\approx\int_{0}^{\infty}D(\omega)\cos(\omega t)\mathrm{d}\omega$ (3)
where the sum has been approximated by an integral. From here, it is clear that the density of states is given by the Fourier transform of $C_{v}(t)$
@@
As seen from the above derivation, there are some key assumptions made about the system when using the VACF method to find the density of states. Specifically, this method is valid for a system meeting the following criteria: the equipartition theorem applies, the system is in thermal equilibrium, and the vibrations are isotropic with random phase. This method also assumes access to the velocities of all the particles.
Given these assumptions, the applicability of this method to granular systems is not guaranteed. However, the VACF method was experimentally applied to a 2D granular packing in Ref. [31] with the authors calling the result the “density of modes” to distinguish it from the thermal density of states. In these experiments, the granular packing was excited with a steady white noise acoustic signal to mimic thermal motion. However, the system demonstrated non-thermal motion, violating a key assumption of the VACF method. In spite of this, the experiments recovered several important features of the density of state, such as Debye scaling in ordered, crystalline packings. It also saw an excess number of low frequency modes compared to Debye scaling in disordered packings and was able to define an $\omega^{*}$ that scaled with $z-z_{c}$ but which scaled only weakly with pressure. These results demonstrate the utility of applying the VACF method to granular systems.
@@
Figure 1: (a) A schematic of the experimental setup. The granular material, shown as the dotted fill, is confined inside a box, shown as the black line surrounding the granular material. Inside the granular material, labeled with a “P” in the figure are the piezoelectric sensors. On top of the granular material is the pressure plate assembly. The pressure plate assembly houses the force sensors that measure the pressure applied to the granular material via the screw which is itself confined to the frame. Vibrations are excited in the granular material when the ball is released from the solenoid and dropped through a tube onto the bottom plate of the pressure plate assembly. The bottom plate of the pressure plate assembly is in direct contact with the top of the granular material. (b) A photo showing the ball dropping apparatus with the red wires composing the solenoid and the aluminum tube below the solenoid. To the right, and from the top down, are as follows: the piezoelectric sensor encased in the ABS half spheres, the bare piezoelectric sensor, a 6 mm grain, a 8 mm grain, and the steel impactor ball.
A key aspect of applying the VACF method to a granular material is to excite the material to create particle velocities that are not normally present in a granular material. In Ref. [31], the granular material was excited with a white noise acoustic signal to mimic thermal motion. This method of exciting the system can be thought of as similar to a damped, driven oscillator where the system is forced to oscillate at the drive frequencies, but the amplitude of these oscillations are attenuated in relation to the natural resonances of the system. That is, the system will attenuate frequencies that are not resonant with the system while minimally attenuating resonant frequencies. This method of exciting the system has an advantage that dissipation in the granular material is counteracted due to the constant white noise drive. However, there are several significant experimental drawbacks related to the complexity of implementing this method. Specifically, a very flat spectrum of white noise must be injected into the system due to the fact that any electrical or mechanical resonances in the driving apparatus will be transmitted to the density of states. Compensating for these resonances requires careful design of electronics and calibration making this a difficult excitation method to apply outside of carefully controlled laboratory conditions. Additionally, since the system is being continuously driven, a relatively large and powerful driving mechanism must be used. An alternative method of exciting the vibrational modes that used the VACF method is reported in Ref. [32] where they used the natural phonon emissions of stick slip events of a granular material under shear to compute the density of modes. This method eliminated the above mentioned experimental difficulties in the driving method; however, this method is not applicable to static granular packings due to the lack of shear.
Drawing from the work of Refs. [31, 32], we propose a method to excite vibrational modes in a granular material for the purpose of measuring the granular density of modes with the VACF method. Our method excites the vibrational modes using a small impact on top of the granular material. This impact will create a pulse containing a wide spectrum of frequencies. The range of frequencies in a pulse is inversely proportional to the width of the pulse, with an infinitesimal pulse having a flat spectrum. This wide injection of frequencies will allow the granular material to vibrate more strongly at its natural resonances and attenuate signals that are not resonant with the natural modes of the system.
This impact method of excitation is experimentally simple; it does not require complex electronics, large shakers, or the need to compensate for apparatus resonances. In addition to being experimentally simple, an impact has a particular intrinsic advantage. Namely, while the impactor is striking the granular material, the system can also be thought of as a damped, driven oscillator; however, since the impact event is very short, transients will be important to the response. For damped, driven oscillators, the transient terms oscillate at the natural frequency of the system and not the driving frequency, which is exactly what is needed for our experiments. Then, after the impact event, the driving force is no longer present and the system behaves more like a damped oscillator, essentially ringing at the natural frequencies of the material. This allowance of transients is in contrast to previously used continuous excitation methods [31] where the system is forced to oscillate at all modes and then attenuate modes that are not resonant with the system. The primary disadvantage of this impact technique is also due to the transient nature of the excitation, in that, it may be difficult to overcome dissipation as the vibrations travel through the material.
In this paper, we experimentally measure the granular density of modes of a real three-dimensional granular material utilizing the VACF method with an impact event to excite the vibrational modes. Our results recover several previously discussed features of the granular density of states, such as an excess of low frequency modes at low pressure. Additionally, this work adds another experimental system to the relatively small number of systems for which there is data on the granular density of modes.
@@ II Experimental Methods
For our experiment, we investigate a granular material confined within a box closed on five sides and open at the top, see Figure 1. The granular material inside the box has dimensions of $25\times 25\times 25$ cm3 and is composed of one of the following three possible grain types: i) a monodisperse packing of 6 mm spherical grains, ii) a monodisperse packing of 8 mm spherical grains, iii) a 50-50 bidisperse mixture by volume of 8 mm and 6 mm spherical grains. For both grain sizes, the grains are composed of ABS plastic with a Young’s modulus of $E\approx 2.6$ GPa [34], a Poisson’s ratio of $\nu\approx 0.37$ [35, 36], and individual grain masses of $m_{6mm}=0.20$ g and $m_{8mm}=0.34$ g.
In order to shift the granular material further from the jamming transition, pressure is applied to the granular material via a pressure plate apparatus placed on top of the granular system. This pressure plate apparatus consists of two $25\times 25$ cm2 parallel plates with three force sensors (TE Connectivity Measurement Specialties, FX19 Series sensors) placed in between the plates to allow us to measure the total force applied to the system. Pressure is then applied to the pressure plate apparatus via a screw built into a frame that encircles the grain box, see Figure 1. This apparatus for applying pressure to the granular material is similar to a cider press for pressing apples into juice, where the grains are the “apples” in this analogy. For our experiments, the lowest pressure is achieved with no force applied from the screw; then, we apply the following pressures to the system by tightening the screw and compressing the granular material from the top: 0 kPa, 3.6 kPa, 7.1 kPa, 14.2 kPa. Since the sensors are buried in approximately the middle of the granular packing, we add the weight of the grains above the sensors (50 N) for a revised pressure range of 0.8 kPa, 4.4 kPa, 7.9 kPa, 15 kPa. For a more intuitive understanding of these pressure ranges, the pressures can be expressed as fractions of the Young’s modulus of the grains as follows: $0.3~{}\mu E$, $1.7~{}\mu E$, $3.0~{}\mu E$, $5.8~{}\mu E$, where the unit $\mu E=10^{-6}E$.
@@
II.1 Piezoelectric sensors
We use five piezoelectric sensors embedded in the granular material to measure the acoustic vibrations of the system caused by the impact of a steel ball being dropped on the granular material. The particular piezoelectric sensors used are Murata 7BB-12-9 buzzers. While these crystals are marketed as buzzers, they work equally well as detectors due to the symmetric nature of the piezoelectric effect: that is, a strain on a piezoelectric crystal will produce charge separation and, inversely, a charge difference across a piezoelectric crystal will produce a strain. The piezoelectric sensors are 12 mm in total diameter with the crystal being 9 mm in diameter. For each sensor, we glue two ABS plastic half spheres on either side in order to turn the sensor into a 12 mm diameter ABS sphere with a piezoelectric sensor running through its equator, see Figure 1. Encasing the sensors in ABS spheres allows the sensor to better blend in with the rest of the material which is also composed of ABS plastic spheres. This helps to ensure that the sensors are vibrating in as close a manner to the bulk granular material as possible.
Figure 4: The unfiltered spectrum of a piezoelectric affixed to the top of the impact plate showing both the injected frequencies from the impact as well as the noise when there is no impact. The noise spectrum shows the classic $\frac{1}{f}$ at low frequency followed by broadband noise [37], and we fit the noise spectrum using this form for reference and later use. The noise spectrum also shows several narrow spikes which are harmonics of the 60 Hz mains. These harmonics of 60 Hz are filtered from the later dataset using a digital notch filter with a bandwidth on the order of 1 Hz.
If these sensors are simply placed into the granular material without any additional circuitry, a voltage could be measured in some proportion to acoustic vibrations. However, this signal would be very weak, possibly not even exceeding the noise threshold. It is therefore necessary to connect the piezoelectric sensors to pre-processing circuitry to amplify the signal and reduce the noise level. The circuit that we developed (shown in Figure 2) to accomplish this task consists of two charge amplifiers [38], one connected to each side of the crystal. These charge amplifiers convert the charge on each side of the crystal into a voltage. These two voltages are then fed through a differential amplifier for both amplification and common mode noise rejection. This circuit yields high sensitivity and low noise, ideal for measuring small vibrations. Additionally, the output voltage of the circuit, $V_{out}$ can be converted into the charge, $q$, generated by the piezoelectric with the following relation $q=\frac{V_{out}}{2GC_{f}}$, where $G$ is the gain of the differential amplifier and $C_{f}$ is the feedback capacitor. Figure 3 shows the output of a typical piezoelectric sensor.
II.2 Phonon excitation
In order to excite the granular material to vibrate at its natural modes, we drop a 1.05 g, 6.35 mm diameter, steel ball, see Figure 1, on top of the system, thereby injecting an acoustic pulse into the system. In order to produce a consistent pulse across our different packings and pressures, the following apparatus was constructed to drop the pulse-producing steel ball on the granular material. The pulse producing apparatus (labeled “ball” in Figure 1) consists of a 5.0 cm tall hollow aluminum tube on top of which is a solenoid. Inside and at the top of the tube, but below the solenoid, the steel ball is placed. The steel ball is held in place by the solenoid’s magnetic field and can be dropped when the solenoid is turned off. This setup allows for a consistent kinetic energy of the ball (0.5 mJ) as it strikes the top plate of the system, as well as provides a means of coordinating the time at which the ball is dropped with the data acquisition of the embedded piezoelectric sensors. Additionally, by striking the rigid plate of the pressure plate apparatus, the energy of the pulse is more evenly distributed across the granular material rather than being concentrated in a particular region of the system.
In order to excite a broad spectrum of frequencies with this impact technique, the impact event should be as short as possible. To design a system that facilities a short impact time, we imagine the impactor colliding with the impactor plate as a mass colliding with a spring, then the period of that spring’s oscillation would correspond to the duration of the collision. Using this model for the impact time, a short impact can be accomplished with a light, but hard and stiff impactor. If higher energy is needed from the impactor, it is best accomplished by increasing the speed of the impactor rather than its mass so that the impact time remains small. Finally, so long as the energy of the impactor does not cause the grains to permanently change their position, the energy of the impactor should not influence the $D(\omega)$ of an individual packing, provided there is enough energy to excite the same number of modes with different energies. However, if data from multiple packing are averaged together, as is done in this study, the energy of the impactor should be kept very consistent as packings with higher energy impacts will be weighted more heavily in Equation 1.
Keeping all of these design principles in mind, we chose our impactor to be made of a hardened steel ball bearing with a mass equal to only about five of the 6 mm grains or three of the 8 mm grains. We chose the drop height, and thus the speed of the impactor, by dropping the impactor from the highest height that did not cause the piezoelectric sensor voltages to be clipped.
We measure the range of frequencies produced with our impact setup as well as the electrical noise of our circuit by attaching a piezoelectric sensor to the top of the impactor plate and dropping the impactor onto the plate; the results of this test are shown in Figure 4. Since the piezoelectric sensor is in close proximity to the impactor, it has a larger amplitude than the piezoelectrics in the granular material, and we therefore omit the first five bounces of the impactor as the signal from these bounces is clipped by our circuit. Omitting these first bounces simply means that the amplitude of the injected frequencies is greater than what is shown in Figure 4 and the curve in Figure 4 represents a floor on the amplitude of the injected frequencies. Additionally, the impactor plate is resting on the granular material and the structure seen in the injected signal is likely due to the granular material. Finally, we see that the injected frequencies are above the noise threshold for the frequencies we investigate, demonstrating that our setup is exciting a broad band of frequencies through the measurement limit of 25 kHz.